A very long, straight wire has charge per unit len
Answer #1

Here ,

string charge density , lamda = 2.3 *10^-10 C

let the distance from the wire is d

as the electric field due to wire is given as

E = 2*k *lamda/(d)

8 = 2 * 9*10^9 * 2.3 *10^-10/d

solving for d

d = 0.5175 m

the distance from the wire is 0.5175 m